# Data-based control of Logical Networks

### v1

@@ Abstract
In recent years, data-driven approaches have become increasingly pervasive across all areas of control engineering. However, the applications of data-based techniques to Boolean Control Networks (BCNs) are still very limited. In this paper we aim to fill this gap, by exploring the possibility of solving three fundamental control problems, i.e., state feedback stabilization, safe control and output regulation, for a BCN, leveraging only a limited amount of data generated by the network, without knowing or identifying its model.
@@
Boolean control networks; feedback stabilization; safe control; output regulation.
@@ 1 Introduction
Boolean Networks (BNs) are discrete-time (autonomous) dynamical systems whose describing variables are constrained to take values in the Boolean set. When their evolution is controlled by an external signal (an input), whose entries in turn take Boolean values, they are called Boolean Control Networks (BCNs). The interest in this class of logical systems originated in the 1960s, when Stuart Kauffman introduced them to model and analyze gene regulatory networks Kauffman (1969), and even if this may still be regarded as their most successful application area, they have been adopted in a wide range of fields, including systems biology, computer science, game theory, smart cities and network theory (see, e.g., Bornholdt (2008); Cheng (2014); Green et al. (2007); Kabir et al. (2014)).
@@
The algebraic approach to BNs and BCNs, proposed by Cheng et al. Cheng (2009); Cheng & Qi (2010a, b); Cheng et al. (2011), and deeply relying on the concept of semi-tensor product (STP) Cheng (2001), has made it possible to represent BNs and BCNs by means of discrete-time state space models in which the state, input and output are logical vectors, and the describing matrices are logical matrices, in turn. In this way, concepts and tools developed for linear time-invariant state space models have been adjusted to formalize and solve control problems for BNs and BCNs (see, e.g., Fornasini & Valcher (2014); Li et al. (2013)).
@@
Despite the simplifying “on/off" (“active/inactive", “high/low") logic at the basis of BCNs, the size and complexity of the physical systems they model (e.g., gene regulatory networks) make it difficult to derive an accurate model of the network, as well as to obtain it by means of identification techniques. On the other hand, in recent times direct data-driven methods have gained interest in the control community, given their attractive feature of bypassing the identification phase, to directly design the problem solution based on raw data. Yet, the use of partial data to solve control problems for logical networks is still at an early stage, and to the best of our knowledge the first and only attempt in this direction has been made in Leifeld et al. (2018), where a one-step data-driven approach to solve the output stabilization problem is proposed. More specifically, the authors in Leifeld et al. (2018) first reconstruct from the available data the output prediction matrix that maps the current input/output pair to the set of all possible next outputs, and then, based on this map, design an output feedback law to stabilize the output of the BCN to a desired value. However, in general, the set of all possible next outputs contains more than one element and hence the output prediction is possibly non deterministic. This uncertainty is handled by resorting to Probabilistic Boolean Networks (PBNs) to compensate for the fact that data are not sufficient to identify the whole BCN and uniquely solve the stabilization problem.
@@
In this paper, we adopt a different viewpoint and exploit the concept of informativity introduced by van Waarde, Eising, and coauthors in van Waarde et al. (2020) (see, also, Waarde et al. (2023)), to address the limitations arising from having only a limited amount of data available, without the need to resort to PBNs. Indeed, the informativity framework is based on the idea that, when data are not informative enough to uniquely identify the system, the best that we can do is imposing that all the systems compatible with the collected data (among which there is surely also our target system) possess the desired property or behave in a certain way. Therefore, in this work we solve three fundamental control problems, namely stabilization, safe control and output regulation of a BCN by state feedback, by solving them for all the logical systems compatible with the available data. This automatically guarantees that the problem is solved for the actual BCN, even under partial information. More specifically, we provide necessary and sufficient conditions on the collected data to stabilize a BCN to a certain state, to drive its evolution towards a safe set, and to set its output to a desired value, by means of a state feedback. Moreover, for each of these problems, we propose a procedure to design such a feedback matrix based on the collected data.
The paper is organized as follows. Section 2 contains some preliminary notions and results about BNs and BCNs. In Section 3 we formalize and solve the data-driven feedback stabilization problem for a BCN. The safe control problem is investigated in Section 4. In Section 5 we state and solve the data-driven output regulation problem by state feedback for a BCN. Section 6 concludes the paper.
This paper extends a manuscript recently submitted to CDC 2025. Compared to the conference version the main novel contribution is represented by Section 4.
@@
Notation. Given two nonnegative integers $k$ and $n$, with $k\leq n$, we denote by $[k,n]$ the set of integers $\{k,k+1,\dots,n\}$. We consider Boolean vectors and matrices whose entries take values in $\mathcal{B}\triangleq\{0,1\}$, equipped with the usual logical operations, namely sum (OR) $\vee$, product (AND) $\land$ and negation $\neg$. We denote by $\delta_{k}^{i}$ the $k$-dimensional $i$-th canonical vector, by $\mathcal{L}_{k}$ the set of all $k$-dimensional canonical vectors, and by $\mathcal{L}_{k\times q}\subset\mathcal{B}^{k\times q}$ the set of all $k\times q$ logical matrices, namely matrices whose $q$ columns are canonical vectors of size $k$. Any matrix $L\in\mathcal{L}_{k\times q}$ can be described as $L=[\delta_{k}^{i_{1}}\ \delta_{k}^{i_{2}}\ \dots\ \delta_{k}^{i_{q}}]$, for some indices $i_{1},i_{2},\dots,i_{q}\in[1,k]$. $I_{n}$ denotes the $n$-dimensional identity matrix. The $k$-dimensional vector with all zero entries is denoted by $\mathbbold{0}_{k}$, and the $k\times l$ matrix with all zero entries by $\mathbbold{0}_{k\times l}$. The $(i,j)$-th entry of a matrix $M$ is denoted by $[M]_{ij}$, and the $j$-th entry of a vector ${\bf v}$ by $[{\bf v}]_{j}$. The $j$-th column of matrix $M$ is denoted by ${\rm col}_{j}(M)$. We denote by $\odot$ the Hadamard (or entry-wise) product. A Boolean (in particular, a logical) square matrix $L$ of size $k$ is reducible if there exists a permutation matrix $P\in\mathcal{L}_{k\times k}$ such that
$P^{\top}LP=\begin{bmatrix}L_{11}&L_{12}\cr 0&L_{22}\end{bmatrix},$
@@
$A\ltimes B\triangleq(A\otimes I_{l/n})(B\otimes I_{l/p}),\quad l\triangleq{\rm l.c.m.}\{n,p\}.$
For the properties of the semi-tensor product we refer to Cheng et al. (2011). By exploiting the semi-tensor product, it is possible to extend the previously mentioned bijective correspondence to a bijective correspondence between $\mathcal{B}^{n}$ and $\mathcal{L}_{2^{n}}$, as follows. Given a vector $X=[X_{1}\ X_{2}\ \dots\ X_{n}]^{\top}\in\mathcal{B}^{n}$, set
@@
It is possible to associate with a Boolean (in particular, a logical) matrix $L\in{\mathcal{B}}^{k\times k}$ a directed graph (or, digraph), $\mathcal{D}(L)=(\mathcal{V},\mathcal{E})$, where ${\mathcal{V}}=[1,k]$ is the set of nodes and ${\mathcal{E}}$ is the set of edges. An edge $(j,l)$, from $j$ to $l$, belongs to $\mathcal{E}$ if and only if $[L]_{lj}=1$. A sequence $j_{1}\to j_{2}\to\dots\to j_{r}\to j_{r+1}$ in $\mathcal{D}(L)$ is a path of length $r$ from $j_{1}$ to $j_{r+1}$ if $(j_{1},j_{2}),\dots,(j_{r},j_{r+1})$ are edges in $\mathcal{E}$. A closed path is a cycle. A node $j^{*}\in{\mathcal{V}}$ is said to be globally reachable if there exists a path to $j^{*}$ from any other node in the network. A digraph $\mathcal{D}(L)$ is strongly connected if all its nodes are globally reachable, and this is the case if and only if $L$ is irreducible.
2 Boolean Control Networks: Preliminaries
@@ A Boolean control network (BCN) is a logical system described by the following equations
where ${\bf x}(t)\in\mathcal{L}_{N}$ and $L\in{\mathcal{L}}_{N\times N}$, with $N=2^{n}$. Similarly, the algebraic representation of the BCN (1) is
@@
where ${\bf x}(t)\in\mathcal{L}_{N}$, ${\bf u}(t)\in\mathcal{L}_{M}$, ${\bf y}(t)\in\mathcal{L}_{P}$, $L\in\mathcal{L}_{N\times NM}$ and $H\in{\mathcal{L}}_{P\times N}$, with $N=2^{n}$, $M=2^{m}$ and $P=2^{p}$. It is worth remarking that all the subsequent analysis holds for generic nonnegative integer values of $N,M$ and $P$, not necessarily powers of $2$. The matrix $L\in\mathcal{L}_{N\times NM}$, whose columns are canonical vectors of size $N$, can be divided into $M$ square blocks of size $N$ as follows
$L=\left[\begin{array}[]{c|c|c|c}L_{1}&L_{2}&\dots&L_{M}\end{array}\right],$ (5)
@@
Cheng et al. (2011) Given a BCN as in (4a), we say that ${\bf x}_{f}=\delta_{N}^{i}$ is reachable from ${\bf x}_{0}=\delta_{N}^{j}$ if there exist $\tau\in\mathbb{Z}_{+}$ and an input ${\bf u}(t),t\in[0,\tau-1],$ that leads the state trajectory from ${\bf x}(0)={\bf x}_{0}$ to ${\bf x}(\tau)={\bf x}_{f}$. The whole BCN is reachable if the previous condition holds for every ${\bf x}_{0}$ and ${\bf x}_{f}$ in ${\mathcal{L}}_{N}$.
We can characterize the reachability of a single state and of the whole network Cheng et al. (2011) by resorting to the matrix

### v2

@@ Abstract
In recent years, data-driven approaches have become increasingly pervasive across all areas of control engineering. However, the applications of data-based techniques to Boolean control networks (BCNs) are still very limited. In this paper we aim to fill this gap, by exploring the possibility of evaluating some basic features, i.e., reachability and equilibria, and of solving two fundamental control problems, i.e., safe control and output regulation, for a BCN, leveraging only a limited amount of data generated by the network, without knowing or identifying its model.
@@
Boolean control networks; safe control; output regulation.
@@ 1 Introduction
Boolean networks (BNs) are discrete-time (autonomous) dynamical systems whose describing variables are constrained to take values in the Boolean set. When their evolution is controlled by an external signal (an input), whose entries in turn take Boolean values, they are called Boolean control networks (BCNs). The interest in this class of logical systems originated in the 1960s, when Stuart Kauffman introduced them to model and analyze gene regulatory networks Kauffman (1969), and even if this may still be regarded as their most successful application area, they have been adopted in a wide range of fields, including systems biology, computer science, game theory, smart cities and network theory (see, e.g., Bornholdt (2008); Cheng (2014); Green et al. (2007); Kabir et al. (2014); Schreiber & Valcher (2022)).
@@
The algebraic approach to BNs and BCNs, proposed by Cheng et al. Cheng (2009); Cheng & Qi (2010a, b); Cheng et al. (2011), and deeply relying on the concept of semi-tensor product (STP) Cheng (2001), has made it possible to represent BNs and BCNs by means of discrete-time state space models in which the state, input and output are logical vectors, and the describing matrices are logical matrices. In this way, concepts and tools developed for linear time-invariant state space models have been adjusted to formalize and solve control problems for BNs and BCNs (see, e.g., Fornasini & Valcher (2014); Li et al. (2013); Zhang & Zhang (2016)).
@@
Despite the simplifying “on/off" (“active/inactive", “high/low") logic at the basis of BCNs, the size and complexity of the physical systems they model (e.g., gene regulatory networks) make it difficult to derive an accurate model of the network, as well as to obtain it by means of identification techniques. On the other hand, in recent times direct data-driven methods have gained interest in the control community, given their attractive feature of bypassing the identification phase, to directly design the problem solution based on raw data. Yet, the use of partial data to solve control problems for logical networks is still at an early stage. The first attempt in this direction has been made in Leifeld et al. (2018), where a one-step data-driven approach to solve the output stabilization problem is proposed. More specifically, the authors in Leifeld et al. (2018) first reconstruct from the available data the output prediction matrix that maps the current input/output pair to the set of all possible next outputs, and then, based on this map, design an output feedback law to stabilize the output of the BCN to a desired value. However, in general, the set of all possible next outputs contains more than one element and hence the output prediction is possibly non-deterministic. To deal with this uncertainty, the authors draw inspiration from results on the stabilization of Probabilistic Boolean networks (PBNs).
The authors in Li et al. (2025), instead, exploit the informativity framework, introduced by van Waarde and coauthors in van Waarde et al. (2020) (see, also, van Waarde et al. (2023)), to investigate BCN control problems in a data-driven set-up. The informativity framework is based on the idea that, when data are not informative enough to uniquely identify the system, the best we can do is impose that all the systems compatible with the collected data (among which there is surely also our target system) possess the desired property or behave in a certain way. In Li et al. (2025) fundamental system properties such as controllability and stabilizability are investigated, and the problems of state feedback stabilization and guaranteed cost control are addressed for BCNs. For both problems, necessary and sufficient conditions under which the data are informative for control design are provided.
@@
In this paper, we also exploit this approach to address the limitations arising from having only a limited amount of data available. More specifically, we solve two fundamental control problems, namely safe control and output regulation of a BCN via state feedback, by solving them for all logical systems compatible with the available data. This automatically guarantees that the problem is solved for the actual BCN, even under partial information. We provide necessary and sufficient conditions on the collected data to drive the evolution of a target BCN towards a safe set, and to set its output to a desired value, by means of a state feedback. Moreover, for each of these problems, we propose a procedure to design such a feedback matrix based on the collected data. In order to solve these problems, we preliminarily investigate under what conditions the reachability of a set of states and the existence of equilibrium points can be inferred, for all the BCNs compatible with the data, directly from the data themselves, since these notions prove to be instrumental in constructing the solutions.
The paper is organized as follows. In Section 2 we briefly introduce Boolean networks and Boolean control networks. In Section 3 we explain the data collection process and we provide the notion of informativity for identifiability, along with its characterization. Section 4 contains an informativity-based characterization of the reachability property for a BCN. Section 5 reviews the notion of equilibrium points in BCNs and provides a systematic approach to identifying all equilibrium points consistent with the data. In Section 6 we formalize and solve the safe control problem for a BCN using only finite sequences of data. In Section 7 the data-driven output regulation problem by state feedback for a BCN is investigated. Section 8 concludes the paper.
@@
Notation. Given two nonnegative integers $k$ and $n$, with $k\leq n$, we denote by $[k,n]$ the set of integers $\{k,k+1,\dots,n\}$. We consider Boolean vectors and matrices whose entries take values in $\mathcal{B}\triangleq\{0,1\}$, equipped with the usual logical operations, namely sum (OR) $\vee$, product (AND) $\land$, and negation (NOT) $\neg$. We denote by $\delta_{k}^{i}$ the $k$-dimensional $i$-th canonical vector, by $\mathcal{L}_{k}$ the set of all $k$-dimensional canonical vectors, and by $\mathcal{L}_{k\times q}\subset\mathcal{B}^{k\times q}$ the set of all $k\times q$ logical matrices, namely matrices whose $q$ columns are canonical vectors of size $k$. Any matrix $L\in\mathcal{L}_{k\times q}$ can be described as $L=[\delta_{k}^{i_{1}}\ \delta_{k}^{i_{2}}\ \dots\ \delta_{k}^{i_{q}}]$, for some indices $i_{1},i_{2},\dots,i_{q}\in[1,k]$. $I_{n}$ denotes the $n$-dimensional identity matrix. The $k$-dimensional vector with all unitary (all zero) entries is denoted by $\mathbbold{1}_{k}$ (by $\mathbbold{0}_{k}$), and the $k\times l$ matrix with all zero entries by $\mathbbold{0}_{k\times l}$. When clear from the context or unnecessary, the suffixes are omitted. The $(i,j)$-th entry of a matrix $M$ is denoted by $M]_{ij}$, and the $j$-th entry of a vector $\bf v$} by ${\bf v}]_{j}$. The $j$-th column of matrix $M$ is denoted by ${\rm col}_{j}(M)$. We denote by $\odot$ the Hadamard (or entry-wise) product. Given two Boolean matrices $M_{1}$ and $M_{2}$ of the same size, condition $M_{1}\leq M_{2}$ means that $M_{1}]_{ij}\leq[M_{2}]_{ij}$ for every $i,j$. Given a vector ${\bf v}\in{\mathbb{R}}^{k}$, we denote by ${\rm diag}({\bf v}^{\top})$ the $k\times k$ diagonal matrix whose $(i,i)$-th entry is ${\bf v}]_{i}$. A Boolean (in particular, a logical) square matrix $L$ of size $k$ is reducible if there exists a permutation matrix $P\in\mathcal{L}_{k\times k}$ such that
$P^{\top}LP=\begin{bmatrix}L_{11}&L_{12}\cr\mathbbold{0}&L_{22}\end{bmatrix}$,
@@
$A\ltimes B\triangleq(A\otimes I_{l/n})(B\otimes I_{l/p}),\quad l\triangleq{\rm l.c.m.}\{n,p\},$
where the symbol $\otimes$ denotes the Kronecker product. For the properties of the semi-tensor product, we refer to Cheng et al. (2011).
Given two matrices $C\in{\mathcal{B}}^{m\times n}$ and $D\in{\mathcal{B}}^{p\times n}$ with the same number of columns, we denote by * their Khatri-Rao product, which is defined as a column-wise semi-tensor product, or, equivalently, a column-wise Kronecker product, i.e.,
$C*D$ $=$ $\left[\begin{array}[]{c|c|c}{\rm col}_{1}(C)\ltimes{\rm col}_{1}(D)&\dots&{\rm col}_{n}(C)\ltimes{\rm col}_{n}(D)\end{array}\right]$
$=$ $\left[\begin{array}[]{c|c|c}{\rm col}_{1}(C)\otimes{\rm col}_{1}(D)&\dots&{\rm col}_{n}(C)\otimes{\rm col}_{n}(D)\end{array}\right].$
By exploiting the semi-tensor product, it is possible to extend the previously mentioned bijective correspondence to a bijective correspondence between $\mathcal{B}^{n}$ and $\mathcal{L}_{2^{n}}$, as follows. Given a vector $X=[X_{1}\ X_{2}\ \dots\ X_{n}]^{\top}\in\mathcal{B}^{n}$, set
@@
It is possible to associate with a Boolean (in particular, a logical) matrix $L\in{\mathcal{B}}^{k\times k}$ a directed graph (or, digraph), $\mathcal{D}(L)=(\mathcal{V},\mathcal{E})$, where ${\mathcal{V}}=[1,k]$ is the set of nodes and $\mathcal{E}$} is the set of edges. An edge $(j,l)$, from $j$ to $l$, belongs to $\mathcal{E}$ if and only if $L]_{lj}=1$. A sequence $j_{1}\to j_{2}\to\dots\to j_{r}\to j_{r+1}$ in $\mathcal{D}(L)$ is a path of length $r$ from $j_{1}$ to $j_{r+1}$ if $(j_{1},j_{2}),\dots,(j_{r},j_{r+1})$ are edges in $\mathcal{E}$. A closed path is a cycle. A node $j^{*}\in{\mathcal{V}}$ is said to be globally reachable if there exists a path to $j^{*}$ from any other node in the network. A digraph $\mathcal{D}(L)$ is strongly connected if all its nodes are globally reachable, and this is the case if and only if $L$ is irreducible.
2 Introduction to Boolean control networks
@@ A Boolean control network (BCN) is a logical system described by the following equations
where ${\bf x}(t)\in\mathcal{L}_{N}$ and $L\in{\mathcal{L}}_{N\times N}$, with $N\triangleq 2^{n}$. Similarly, the algebraic representation of the BCN (3) is
@@
where ${\bf x}(t)\in\mathcal{L}_{N}$, ${\bf u}(t)\in\mathcal{L}_{M}$, ${\bf y}(t)\in\mathcal{L}_{P}$, $L\in\mathcal{L}_{N\times NM}$ and $H\in{\mathcal{L}}_{P\times N}$, with $N\triangleq 2^{n}$, $M\triangleq 2^{m}$ and $P\triangleq 2^{p}$. It is worth remarking that model (6) can be used to represent any state-space model whose state, input and output variables take values in sets of finite cardinalities $N$, $M$ and $P$, respectively, and hence all the subsequent analysis holds for generic nonnegative integer values of $N$, $M$ and $P$, not necessarily powers of $2$. The matrix $L\in\mathcal{L}_{N\times NM}$, whose columns are canonical vectors of size $N$, can be divided into $M$ square blocks of size $N$ as follows
$L=\begin{bmatrix}L_{1}&\vline&L_{2}&\vline&\dots&\vline&L_{M}\end{bmatrix}$, (7)
@@
3 Data collection process and informativity for identifiability of BCNs
We assume to have performed some offline experiments [^1] It is worth noticing that the number and the lengths of the experiments will play no role in the following analysis. However, a higher number of experiments may allow to obtain information that could not be collected in a single one., say $r\geq 1$, during which we have collected state/input/output data from the BCN (6) on finite time intervals $0,T_{i}],T_{i}\in\mathbb{Z}_{+},i\in[1,r]$. We define the vector sequences ${\bf x}_{d}^{i}\triangleq\{{\bf x}_{d}^{i}(t)\}_{t=0}^{T_{i}}$, ${\bf u}_{d}^{i}\triangleq\{{\bf u}_{d}^{i}(t)\}_{t=0}^{T_{i}-1},$ and ${\bf y}_{d}^{i}\triangleq\{{\bf y}_{d}^{i}(t)\}_{t=0}^{T_{i}-1},i\in[1,r]$, and accordingly ${\bf x}_{d}\triangleq\{{\bf x}_{d}^{i}\}_{i=1}^{r}$, ${\bf u}_{d}\triangleq\{{\bf u}_{d}^{i}\}_{i=1}^{r}$ and ${\bf y}_{d}\triangleq\{{\bf y}_{d}^{i}\}_{i=1}^{r}$. We rearrange the data collected during the $r$ experiments into the following logical matrices with $T\triangleq\sum_{i=1}^{r}{T_{i}}$ columns:
$X_{p}$ $\triangleq$ $\left[\begin{array}[]{ccc|ccc}{\bf x}_{d}^{1}(0)&\dots&{\bf x}_{d}^{1}(T_{1}-1)&{\bf x}_{d}^{2}(0)&\dots&{\bf x}_{d}^{r}(T_{r}-1)\end{array}\right],$
$X_{f}$ $\triangleq$ $\left[\begin{array}[]{ccc|ccc}{\bf x}_{d}^{1}(1)&\dots&{\bf x}_{d}^{1}(T_{1})&{\bf x}_{d}^{2}(1)&\dots&{\bf x}_{d}^{r}(T_{r})\end{array}\right],$
$U_{p}$ $\triangleq$ $\left[\begin{array}[]{ccc|ccc}{\bf u}_{d}^{1}(0)&\dots&{\bf u}_{d}^{1}(T_{1}-1)&{\bf u}_{d}^{2}(0)&\dots&{\bf u}_{d}^{r}(T_{r}-1)\end{array}\right],$
$Y_{p}$ $\triangleq$ $\left[\begin{array}[]{ccc|ccc}{\bf y}_{d}^{1}(0)&\dots&{\bf y}_{d}^{1}(T_{1}-1)&{\bf y}_{d}^{2}(0)&\dots&{\bf y}_{d}^{r}(T_{r}-1)\end{array}\right],$
where the subscripts $p$ and $f$ stand for past and future, respectively.
As every BCN (6) bijectively corresponds to its describing matrices, we define the set of Boolean control networks (6) compatible with the data $({\bf x}_{d}$, ${\bf u}_{d},{\bf y}_{d})$ as
$\!\!\!\!\mathcal{B}_{d}\!$ $\triangleq\big\{(\tilde{L},\tilde{H})\in\mathcal{L}_{N\times NM}\times\mathcal{L}_{P\times N}:$
$\!\!\!{\bf x}_{d}^{i}(t+1)=\tilde{L}\ltimes{\bf u}_{d}^{i}(t)\ltimes{\bf x}_{d}^{i}(t),\ {\bf y}_{d}^{i}(t)=\tilde{H}{\bf x}_{d}^{i}(t),$
$\!\!\!\forall t\in[0,T_{i}-1],\forall i\in[1,r]\big\}.\!\!\!$ (13)
Obviously, the pair $(L,H)$ corresponding to the BCN (6) that generated the data belongs to $\mathcal{B}_{d}$. However, ${\mathcal{B}}_{d}$ may also include additional matrix pairs when the available data do not uniquely determine $(L,H)$. To this end, we introduce the notion of informativity for identifiability [^2] As a matter of fact, what we call informativity for identifiability is referred to as informativity for model reconstructability in Definition 3.3 of Li et al. (2025). , that extends Definition 3.3 in Li et al. (2025) to BCNs as in (6), which also include an output equation.
@@ 3 Data collection process and informativity for identifiability of BCNs
The data $({\bf x}_{d},{\bf u}_{d},{\bf y}_{d})$ are called informative for identifiability of the BCN (6) if ${\mathcal{B}}_{d}\equiv\{(L,H)\}$.
Despite the presence of the output equation, the characterization of informativity for identifiability is the same as the one given in Proposition 3.4 of Li et al. (2025) for BCNs described only by (6a). Lemma 2 proves such equivalence and provides an additional equivalent condition.
Lemma 2.
The following facts are equivalent.
  • i)
    The data $({\bf x}_{d}$, ${\bf u}_{d},{\bf y}_{d})$ are informative for identifiability of the BCN (6).
  • ii)
    For every $(i,j)\in[1,M]\times[1,N]$, there exists $k\in[1,T]$ such that
    $\begin{bmatrix}X_{p}\cr U_{p}\end{bmatrix}\delta_{T}^{k}=\begin{bmatrix}\delta_{N}^{j}\cr\delta_{M}^{i}\end{bmatrix}$. (14)
  • iii)
    The logical matrix $U_{p}*X_{p}$ has no zero rows (equivalently, is of full row rank).
Proof.
i) $\Rightarrow$ iii) If i) holds, then the data $({\bf x}_{d}$, ${\bf u}_{d})$ are informative for identifiability of equation (6a), and hence, by Proposition 3.4 in Li et al. (2025), we can claim that iii) holds.
iii) $\Rightarrow$ i) If condition iii) holds, by Proposition 3.4 in Li et al. (2025), we can uniquely identify the matrix $L$. On the other hand, if the matrix $U_{p}*X_{p}$ is of full row rank, then $X_{p}$ is of full row rank, in turn. This means that all possible states appear at least once in $X_{p}$, and consequently also the corresponding output in $Y_{p}$. This allows to uniquely identify also the matrix $H$ in (6b), and hence the data $({\bf x}_{d},{\bf u}_{d},{\bf y}_{d})$ are informative for identifiability of the whole BCN (6).
ii) $\Leftrightarrow$ iii) is straightforward. ∎
Remark 3.
If the data $({\bf x}_{d},{\bf u}_{d},{\bf y}_{d})$ are informative for identifiability and condition (14) holds, then one can identify the $j$-th column of matrix $L_{i}$ as ${\rm col}_{j}(L_{i})=X_{f}\delta_{T}^{k}$. As it can be deduced from Remark 3.5 in Li et al. (2025), one can identify the whole matrix $L$ using the following formula
$L=X_{f}\odot_{\mathcal{B}}(U_{p}*X_{p})^{\top},$
where $\odot_{\mathcal{B}}$ is the Boolean product of matrices (that acts as the regular matrix product with the multiplication replaced by the AND and the addition replaced by the OR). Alternatively, by resorting to the standard matrix product, one can deduce $L$ as follows:
$L=X_{f}(U_{p}*X_{p})^{\top}\left({\rm diag}(\mathbbold{1}_{N}^{\top}X_{f}(U_{p}*X_{p})^{\top})\right)^{-1}.$
Finally, as $X_{p}$ is of full row rank (see the proof of Lemma 2), it is immediate to obtain $H$ as
$H=Y_{p}X_{p}^{\#},$
where $X_{p}^{\#}\triangleq X_{p}^{\top}(X_{p}X_{p}^{\top})^{-1}$ is a right inverse of $X_{p}$.
To make the following analysis meaningful, from now on we assume that the collected data are not informative for identifiability of the considered BCN, namely ${\mathcal{B}}_{d}\supsetneq\{(L,H)\}$, and we investigate under what conditions it is still possible to use them to evaluate certain properties or to solve some control problems. By adhering to the approach initiated in the papers by van Waarde, Eising et al. (see, e.g., van Waarde et al. (2020, 2023)) and first explored for logical networks in (Li et al., 2025, Definition 2.1), when the available data do not allow identifying the BCN that generated them, the best we can do is to understand if a certain property holds or some control problem is solvable for all the BCNs in ${\mathcal{B}}_{d}$. When so, we will say that the data $({\bf x}_{d},{\bf u}_{d},{\bf y}_{d})$ are informative for the property or for the problem solution.
4 Reachability property
In this section, we provide an informativity-based characterization of the reachability of a set of states for a BCN. Before analyzing the problem in a data-driven context, we first recall the model-based approach to the reachability property.
Definition 4.
Cheng et al. (2011) Given a BCN as in (6a), we say that a state ${\bf x}_{f}=\delta_{N}^{i}$ is reachable from ${\bf x}_{0}=\delta_{N}^{j}$ if there exist $\tau\in\mathbb{Z}_{+}$ and an input ${\bf u}(t),t\in[0,\tau-1],$ that leads the state trajectory from ${\bf x}(0)={\bf x}_{0}$ to ${\bf x}(\tau)={\bf x}_{f}$. A set ${\mathcal{X}}\subset{\mathcal{L}}_{N}$ is reachable from ${\bf x}_{0}$, if there exists ${\bf x}_{f}\in\mathcal{X}$ that is reachable from ${\bf x}_{0}$. A state ${\bf x}_{f}$ (a set $\mathcal{X}$}) is globally reachable if it is reachable from every ${\bf x}_{0}$. The BCN (6a) is reachable if every state ${\bf x}_{f}$ is globally reachable.
The reachability of a single state, of a set $\mathcal{X}$} and of the whole BCN can be characterized (see, e.g., Section 3 in Fornasini & Valcher (2013b) and Chapter 16 in Cheng et al. (2011)) by resorting to the matrix